### v1

# spacing=nonfrench

Abstract
Large Language Models (LLMs) have advanced code generation and software automation, but are fundamentally constrained by limited inference-time context and lack of explicit code structure reasoning. We introduce Kodezi Chronos [51], a next-generation architecture for autonomous code understanding, debugging, and maintenance, designed to operate across ultra-long contexts comprising entire codebases, histories, and documentation—all without fixed window limits.
Kodezi Chronos leverages a multi-level embedding memory engine, combining vector and graph-based indexing with continuous code-aware retrieval. This enables efficient and accurate reasoning over millions of lines of code, supporting repository-scale comprehension, multi-file refactoring, and real-time self-healing actions.
Our evaluation introduces a novel Multi Random Retrieval benchmark, specifically tailored to the software engineering domain. Unlike classical retrieval benchmarks, this method requires the model to resolve arbitrarily distant and obfuscated associations across code artifacts, simulating realistic tasks such as variable tracing, dependency migration, and semantic bug localization. Chronos outperforms prior LLMs and code models—demonstrating a 23% improvement in real-world bug detection and reducing debugging cycles by up to 40% compared to traditional sequence-based approaches.
By natively interfacing with IDEs and CI/CD workflows, Chronos enables seamless, autonomous software maintenance, elevating code reliability and productivity while reducing manual effort. These results mark a critical advance toward self-sustaining, continuously optimized software ecosystems.
I INTRODUCTION
Recent advancements in large language models (LLMs) have transformed code generation, review, and reasoning tasks [1, 2, 10]. However, debugging—the most time-consuming and critical aspect of software development—remains largely unsolved. While tools like GitHub Copilot [11], Cursor, and Claude excel at code completion [12], they fundamentally misunderstand debugging as a multi-faceted, context-heavy process that spans entire repositories, historical commits, CI/CD logs, and runtime behaviors. Production debugging requires reasoning across files separated by thousands of lines [13], understanding temporal code evolution, and correlating seemingly unrelated symptoms to root causes buried deep in dependency chains [14].
Current code assistants fail at debugging for three critical reasons: (1) they are trained primarily on code completion tasks, not debugging workflows [15]; (2) they lack persistent memory of past bugs, fixes, and codebase-specific patterns [16]; and (3) their context windows, even when extended to 100K+ tokens, cannot capture the full debugging context needed for complex, multi-file issues. Recent studies show that even state-of-the-art models like GPT-4, Claude-3, and Gemini-1.5 achieve less than 10% success rates on real-world debugging benchmarks [17, 18], often proposing superficial fixes that fail validation or introduce new regressions [19].
Kodezi Chronos represents a paradigm shift: the first debugging language model developed by Kodezi [51], specifically designed, trained, and optimized for autonomous bug detection, root cause analysis, and validated fix generation. Unlike code completion models that generate syntactically correct but often semantically flawed suggestions, Chronos operates through a continuous debugging loop—proposing fixes, running tests, analyzing failures, and iteratively refining solutions until validation succeeds. Built on a novel architecture combining persistent debug memory, multi-source retrieval (code, logs, traces, PRs), and execution sandboxing, Chronos achieves what no existing model can: true autonomous debugging at repository scale.
Chronos is designed for seamless integration with modern development stacks: it operates as an embedded “AI CTO” within CI/CD pipelines, IDEs, and project management tools. Its proactive, event-driven workflow ensures that debugging, documentation generation, refactoring, and even preventative maintenance actions occur autonomously, guided by deep repository memory rather than manual prompts or brittle heuristics.
To rigorously evaluate Chronos’s unique capabilities, we move beyond traditional benchmarks and propose a multi-step, random retrieval evaluation reflecting the authentic complexities of code search, dependency resolution, and semantic bug localization at scale. Empirically, Chronos demonstrates state-of-the-art results across industry-standard metrics and realistic maintenance tasks, reducing debugging times and increasing project resilience.
This paper presents the architecture, memory system, retrieval mechanism, evaluation methodology, and experimental results that establish Kodezi Chronos as the new frontier for autonomous, ultra-contextual codebase management.
II RELATED WORK AND LIMITATIONS OF TRADITIONAL APPROACHES
The emergence of large-scale neural models for source code processing—such as CodeBERT [46], GraphCodeBERT [4], and CodeT5 [5]—has significantly advanced automatic code synthesis, translation, and review. Many of these models, as well as massive LLMs like GPT-3 and GPT-4 [1, 47], are pre-trained on billions of lines of code paired with natural language, learning rich semantic representations for many programming languages.
Despite impressive gains on benchmark tasks, these approaches are fundamentally bottlenecked by attention-based architectures and fixed-size input windows, typically constraining context to tens of thousands of tokens. Practical developer workflows, by contrast, require tools to reason across entire repositories: files, modules, historic commits, documentation, and evolving build systems, far outstripping standard context limits.
State-of-the-art retrieval-augmented generation (RAG) methods extend practical context by embedding and recalling external documents [8], but remain mostly limited to chunked passage retrieval and lack true compositional, dependency-aware memory. Window expansion techniques—such as models with 100k tokens or more [7]—incur prohibitive compute/memory costs and suffer from diluted attention, leading to information loss and degraded performance as codebase size increases.
Existing benchmarks for long-context reasoning, like the ”Needle in a Haystack” pattern, commonly involve embedding a salient but easily differentiable clue in a sea of noise or repetition. Such evaluations unintentionally rely on the model’s ability to match explicit tokens or unusual patterns, providing only a partial test of true retrieval and reasoning ability. In real-world maintenance, however, code elements are frequently unmarked, distributed, and semantically interdependent, requiring the AI not just to ”find,” but to infer context, resolve relationships, and reason about impact [45].
Other recent work explores graph neural networks (GNNs) for modeling explicit data flow or control flow graphs [9, 6]. While GNNs can capture structural properties of code, they are rarely coupled with ultra-long context LLMs, and lack the continuous learning, memory updating, and rapid recall required for live autonomous maintenance.
Kodezi Chronos is motivated by these challenges: By combining continuous graph-aware indexing, dynamic embedding updates, and reasoning-optimized memory retrieval, Chronos transcends traditional limitations and enables truly repository-scale, real-time software comprehension and intervention.
III ARCHITECTURE OF KODEZI CHRONOS
III-A The Output-Heavy Nature of Debugging
Despite the industry focus on ever-larger context windows (128K, 200K, 1M+ tokens), debugging presents a fundamentally different challenge: it is inherently output-heavy rather than input-heavy. This asymmetry has profound implications for model design and optimization.
III-A1 Input vs Output Token Distribution
What models typically see (input):
  • Error stack traces: 200-500 tokens
  • Relevant source code: 1K-4K tokens
  • Test failures and logs: 500-2K tokens
  • Prior fix attempts: 500-1K tokens
  • Total input: Often ¡10K tokens for most real-world debugging tasks
What models must produce (output):
  • Multi-file bug fixes: 500-1,500 tokens
  • Root cause explanations: 300-600 tokens
  • Updated unit tests: 400-800 tokens
  • Commit messages/PR summaries: 150-300 tokens
  • Documentation updates: 200-400 tokens
  • Total output: Typically 2,000-4,000 tokens per debugging session
III-A2 Why Output Quality Trumps Input Size
TABLE I: Input vs output characteristics in debugging tasks.
Aspect Input Context Output Generation
Nature Sparse, localized Dense, structured
Cost Impact Sublinear with retrieval Linear to exponential
Quality Limiter Retrieval precision Generation accuracy
Success Factor Context relevance Syntactic & semantic correctness
The critical insight: a model with intelligent 8K context that generates robust, test-passing fixes will outperform a 1M-context model that produces syntactically correct but semantically flawed patches.
III-A3 Chronos’s Output-Optimized Architecture
Chronos addresses this asymmetry through several architectural innovations:
  1. 1.
    Debug-Specific Generation Training: Unlike code completion models trained on next-token prediction, Chronos is trained on complete debugging sessions, learning to generate structured fixes, explanations, and tests as cohesive units.
  2. 2.
    Iterative Refinement Loop: Rather than single-shot generation, Chronos validates outputs through execution, using test results to refine patches—ensuring output quality over quantity.
  3. 3.
    Template-Aware Generation: Chronos learns repository-specific patterns for commits, tests, and documentation, reducing output token waste while maintaining consistency.
  4. 4.
    Confidence-Guided Output: The model generates explanations and fallback strategies only when confidence is below threshold, optimizing output token usage.
This output-centric design enables Chronos to achieve 65.3% debugging success despite competitors having 10-100x larger context windows, validating that for debugging, output quality and structure matter more than input capacity.
III-B Core Architecture Overview
Kodezi Chronos is designed as an autonomous memory-driven intelligence layer for code, operating at scales that span entire enterprise repositories, team histories, and auxiliary knowledge sources. Its architecture consists of three core modules: (i) a persistent Memory Engine for continuous graph-based context construction, (ii) an advanced Retriever that dynamically assembles deep semantic context from code and documentation, and (iii) a transformer-based Code Reasoning Model for synthesis, debugging, and orchestration of software changes.
Code, Docs, CI/CD LogsMemory Engine(Embedding + Graph)Multi-CodeAssociation RetrieverReasoning Model& OrchestrationTest ResultsPatches, Changelogs,Test Results
Figure 1: High-level overview of Chronos: Memory-driven embedding and retrieval powering autonomous reasoning and codebase management.
[…]
Figure 2: Token distribution in debugging tasks: Unlike typical LLM applications where input dominates, debugging requires substantial, high-quality output generation.
III-C Memory Engine: Repository-Scale Embedding and Indexing
The Memory Engine is responsible for ingesting, encoding, and maintaining a unified semantic representation of all project files, versioned code, documentation, configuration, historical diffs, test outcomes, and architectural artifacts. Each unit of code (e.g., function level, file, commit) is parsed to extract both sequence and structural cues—using ASTs, dependency graphs, and metadata links.
These elements are then projected into a high-dimensional vector space via context-aware, code-specific encoders trained on multiple programming languages and coding paradigms. The Memory Engine stores not just static embeddings but also maintains an evolving graph database, where nodes represent code elements and edges denote code relationships (e.g., function calls, module imports, bug-ticket links, commit ancestry).
login.pyauth_utils.pyREADME.mdtest_login.pybug_1234.mdcallscited inlinked bug intested by
Figure 3: Graph-structured memory indexing in Kodezi Chronos: code, documentation, and test elements as nodes, with functional relationships as edges.
This design enables Chronos to efficiently retrieve, traverse, and reason about segments of the codebase that share non-local relationships, even if separated by thousands of lines, multiple files, or extensive revision history.
III-D Achieving Unlimited Context Through Smart Retrieval
Traditional LLMs are fundamentally constrained by attention complexity and memory limitations. Even models claiming ”unlimited” context achieve this through sliding windows or hierarchical attention that loses critical debugging information. Chronos implements true unlimited context through:
  • Hierarchical Code Embeddings: Multi-level representations from token → statement → function → module → repository
  • Temporal Context Indexing: Every code element tagged with commit history, allowing time-travel debugging
  • Semantic Dependency Graphs: Explicit modeling of import chains, inheritance hierarchies, and data flows
  • Dynamic Context Assembly: At inference, retrieves precisely the code paths relevant to the current bug
This approach enables Chronos to maintain full repository awareness while operating within reasonable computational bounds—a critical requirement for production deployment.
III-E Deep Context Retrieval and Multi-Code Association
Chronos utilizes a novel multi-code compositional retrieval mechanism, powered by AGR, purpose-built for software engineering. Upon each reasoning or generation request—such as ”fix a failing test,” ”refactor authentication logic,” or ”explain API drift”—the Adaptive Retrieval Engine dynamically assembles a tailored context window by:
  • Issuing semantic queries to the Memory Engine that leverage both metric similarity and structural navigation in the code graph, with dynamic depth expansion (k-hop) based on query complexity.
  • Associating multiple code artifacts through typed relationships: e.g., tracing variable definitions across documentation (k=1), implementation (k=2), regression tests (k=2), and historic bug reports (k=3), stopping when confidence exceeds 90% or diminishing returns detected.
  • Dynamically refining the context through intermediate model inferences and confidence scoring, adapting retrieval depth in real-time. Complex debugging queries automatically trigger deeper graph traversal (k=3-5), while simple lookups terminate at k=1-2.
  • Utilizing edge type priorities: implementation edges (weight=1.0), dependency edges (weight=0.8), documentation edges (weight=0.6), ensuring most relevant paths are explored first.
TABLE II: Example multi-code association retrieval: constructing a task-specific context window for a bug fix.
Step Retrieved Entity Relationship
Q1 login.py Direct bug context
Q2 test_login.py Linked test
Q3 settings.py Imported env vars
Q4 bug_1234.md Historical bug doc
Q5 commit_a1b2c3 Last related commit
This approach allows Chronos to reason across arbitrarily distant, compositionally linked code and documentation artifacts—precisely what is needed for complex debugging, cross-module dependencies, or audit trails.
III-F Adaptive Graph-Guided Retrieval (AGR)
Traditional flat retrieval approaches fail to capture the intricate relationships between code artifacts, leading to incomplete context and erroneous fixes. Chronos introduces Adaptive Graph-Guided Retrieval (AGR), a dynamic mechanism that intelligently expands retrieval neighborhoods based on query complexity and confidence thresholds.
III-F1 Iterative Context Expansion
The AGR mechanism operates through iterative k-hop neighbor expansion:
  1. 1.
    Initial Query Analysis: Decompose the debugging request into semantic components and identify seed nodes in the code graph
  2. 2.
    Adaptive Depth Determination: Calculate optimal retrieval depth based on:
    • Query complexity score (0-1)
    • Code artifact density in the neighborhood
    • Historical debugging patterns for similar issues
  3. 3.
    Guided Expansion: Follow typed edges (implementation, dependency, dataflow) to retrieve contextually relevant nodes
  4. 4.
    Confidence-Based Termination: Stop expansion when retrieval confidence exceeds threshold or diminishing returns detected
III-F2 Graph-Guided vs Traditional Planning
Our empirical analysis reveals fundamental differences between traditional LLM planning and AGR-enhanced debugging:
[…]
Figure 4: Traditional LLM planning vs AGR-enhanced debugging: Graph-guided retrieval provides complete context, leading to accurate implementations.
k=1: Initial RetrievalPlanSignalImplTransEx1Ex2Retrieved:S1_next: Output signal(Type: Signal) k=2: Expanded RetrievalPlanSignalImplTransEx1Ex2StateWaitDocTestAdditional Retrieved:State transitions: Wait-¿S1Example: 9’b101000100100Confidence: 92%
Legend: Query k=1 k=2 Not Retrieved
Figure 5: Iterative context expansion in Adaptive Graph-Guided Retrieval: Starting from a query node, the system progressively expands retrieval depth (k-hops) based on confidence thresholds and query complexity.
BugVectorASTGraphHistoryRetrieverRankerSourceTestsFixesLogs
Figure 6: Multi-modal retrieval mechanism in Chronos.
III-G Reasoning, Generation, and Autonomous Orchestration
The transformer-based Chronos Reasoning Model operates directly over the retrieved, multi-source debugging context. Unlike classical code completion models, Chronos:
  • Diagnoses root causes and synthesizes code changes conditioned on project documentation, prior commits, and dependency patterns.
  • Produces stepwise fix plans, code diffs, documentation updates, and regression test suggestions in a unified, automated debugging loop.
  • Orchestrates a full debugging workflow: proposes bug fixes, invokes relevant tests, parses results, iterates on failures, and generates changelogs or PR summaries—all autonomously.
All outputs and feedback streams (test results, reviewer comments, CI/CD events) are fed back into the Memory Engine, enabling lifelong refinement and rapid adaptation to new debugging scenarios.
This cyclic process of context assembly, reasoning, autonomous validation, and memory update is the core of Chronos’s persistent codebase intelligence, enabling self-sustaining and ever-improving debugging at scale.
Autonomous Code/Documentation Generation(Bug Fix Proposal)AutomatedTesting & ValidationMemory UpdateIterative PlanRefinementTest PassedTest FailedReviewer FeedbackCI/CD Events
Figure 7: Chronos debugging feedback loop: Automated bug fix generation, validation, plan refinement, and memory update for continuous autonomous improvement.
IV THE DEBUGGING LANGUAGE MODEL PARADIGM
Kodezi Chronos fundamentally departs from traditional code models by being purpose-built as a debugging language model—the first of its kind. While existing LLMs treat debugging as a code generation problem, Chronos recognizes it as a complex, iterative process requiring specialized capabilities, training, and architecture.
IV-A Core Architectural Layers
Chronos implements a revolutionary 7-layer architecture specifically designed for autonomous debugging:
  1. 1.
    Multi-Source Input Layer: Ingests heterogeneous debugging signals including source code, CI/CD logs, error traces, stack dumps, configuration files, historical PRs, and issue reports. Unlike code models that primarily process source files, Chronos natively understands debugging artifacts.
  2. 2.
    Adaptive Retrieval Engine: Employs AGR (Adaptive Graph-Guided Retrieval) with a hybrid vector-symbolic approach combining:
    • Dynamic k-hop neighbor expansion based on query complexity
    • AST-aware code embeddings that preserve structural relationships
    • Dependency graph indexing for cross-file impact analysis
    • Call hierarchy mapping for execution flow understanding
    • Temporal indexing of code evolution and bug history
    • Confidence-based termination for optimal context assembly
  3. 3.
    Debug-Tuned LLM Core: A transformer architecture specifically fine-tuned on debugging workflows, not just code completion. Training tasks include:
    • Root cause prediction from symptoms
    • Multi-file patch generation
    • Test failure interpretation
    • Regression risk assessment
  4. 4.
    Orchestration Controller: Implements the autonomous debugging loop:
    • Hypothesis generation from error signals
    • Iterative fix refinement based on test results
    • Rollback mechanisms for failed attempts
    • Confidence scoring for proposed solutions
  5. 5.
    Persistent Debug Memory: Maintains long-term knowledge including:
    • Repository-specific bug patterns and fixes
    • Team coding conventions and preferences
    • Historical fix effectiveness metrics
    • Module-level vulnerability profiles
  6. 6.
    Execution Sandbox: Real-time validation environment supporting:
    • Isolated test execution
    • CI/CD pipeline emulation
    • Performance regression detection
    • Security vulnerability scanning
  7. 7.
    Explainability Layer: Generates human-readable outputs:
    • Root cause explanations with evidence chains
    • Fix rationale documentation
    • Automated PR descriptions and commit messages
    • Risk assessment reports
IV-B Training on Debugging Workflows
Unlike models trained primarily on code completion, Chronos’s training regime focuses exclusively on debugging scenarios:
Pre-training Corpus:
  • 15M+ GitHub issues with linked PRs and fix commits
  • 8M+ stack traces paired with resolutions
  • 3M+ CI/CD logs from failed and fixed builds
  • Production debugging sessions from enterprise partners
  • Open-source bug databases (Defects4J, SWE-bench, BugsInPy)
Specialized Fine-tuning Tasks:
  • Chain-of-Cause Reasoning: Teaching the model to trace error propagation through call stacks and dependencies [49]
  • Multi-Modal Bug Understanding: Correlating code, logs, traces, and documentation
  • Iterative Fix Refinement: Learning from failed fix attempts to improve subsequent proposals [50, 48]
  • Cross-Repository Pattern Recognition: Identifying similar bugs across different codebases
IV-C The Autonomous Debugging Loop
Chronos’s debugging loop represents a fundamental innovation over single-shot code generation:
Detect IssueRetrieve ContextPropose FixRun TestsTests Pass?Commit & DeployRefine StrategyUpdate MemoryYesNo
Figure 8: The Chronos autonomous debugging loop: continuous iteration until validation succeeds.
This loop continues autonomously, with each iteration informed by previous attempts and accumulated knowledge, until a validated fix is achieved or human intervention is requested.
V EVALUATION METHODOLOGY
To rigorously assess Kodezi Chronos’s capabilities across realistic debugging and maintenance workflows, we adopt a multi-faceted evaluation strategy that goes beyond conventional sequence completion or shallow retrieval tests.
V-A Benchmarking Datasets and Tasks
Chronos is evaluated on a diverse suite of benchmarks, comprising:
  • Standard Code Generation Benchmarks: HumanEval [2], MBPP [3], and related tasks for basic code synthesis and function-level reasoning.
  • Repository-Scale Debugging Tasks: A curated set of real-world bug reports, failing test scenarios, and codebase refactoring challenges drawn from open-source projects and industry case studies.
  • Multi-Code Association Retrieval: Custom synthetic benchmarks requiring retrieval and reasoning across arbitrarily placed, subtly linked code/documentation artifacts—simulating variable tracing, cross-file bug localization, or dependency update chains.
V-B Multi-Code Reasoning Evaluation Protocol
Unlike traditional benchmarks that target token-level prediction in narrow context, our protocol explicitly:
  • Randomizes the placement of relevant context (bug source, documentation clue, test assertion) across large codebases and histories.
  • Requires Chronos to retrieve, associate, and utilize multi-code context in a compositional manner, solving tasks that demand reasoning over both explicit code relationships (e.g., function calls, imports) and implicit bug/error propagation patterns.
  • Measures both retrieval accuracy (whether Chronos finds all necessary context) and end-to-end task success (whether it can autonomously fix, validate, and document the issue).
V-C Multi Random Retrieval Benchmark
We introduce the Multi Random Retrieval (MRR) benchmark, specifically designed to evaluate debugging-oriented retrieval capabilities:
V-C1 Benchmark Design
The MRR benchmark consists of 5,000 real-world debugging scenarios where:
  1. 1.
    Context Scattering: Relevant debugging information is randomly distributed across 10-50 files
  2. 2.
    Temporal Dispersion: Critical bug context spans 3-12 months of commit history
  3. 3.
    Obfuscated Dependencies: Variable names and function calls are refactored between bug introduction and discovery
  4. 4.
    Multi-Modal Artifacts: Solutions require combining code, tests, logs, and documentation
V-C2 Evaluation Metrics
  • Retrieval Precision@k: Fraction of retrieved artifacts that are relevant to the bug fix
  • Retrieval Recall@k: Fraction of all relevant artifacts successfully retrieved
  • Fix Accuracy: Whether the generated fix passes all tests and doesn’t introduce regressions
  • Context Efficiency: Ratio of used vs retrieved tokens in the final solution
V-C3 Results on MRR Benchmark
TABLE III: Performance on Multi Random Retrieval benchmark, demonstrating Chronos’s superior ability to find and utilize scattered debugging context.
Model Precision@10 Recall@10 Fix Accuracy Context Eff.
GPT-4 + RAG 42.3% 31.7% 8.9% 0.23
Claude-3 + Vector DB 48.1% 36.2% 11.2% 0.28
Gemini-1.5 + Graph 51.7% 41.8% 14.6% 0.31
Kodezi Chronos 89.2% 84.7% 67.3% 0.71
V-D Adaptive Graph-Guided Retrieval Performance
We evaluate the impact of AGR on debugging accuracy across different retrieval depths:
TABLE IV: Performance metrics for different retrieval strategies. Adaptive AGR dynamically selects optimal k based on query complexity.
Retrieval Method k=1 k=2 k=3 k=adaptive Flat
Precision 84.3±2.1% 91.2±1.4% 88.7±1.8% 92.8±1.2% 71.4±3.2%
Recall 72.1±2.8% 86.4±1.9% 89.2±1.6% 90.3±1.5% 68.2±3.5%
F1 Score 77.7±2.4% 88.7±1.6% 88.9±1.7% 91.5±1.3% 69.8±3.3%
Debug Success 58.2±3.1% 72.4±2.3% 71.8±2.4% 87.1±1.8% 23.4±4.1%
Key findings from AGR evaluation:
  • Optimal Depth Varies: Simple bugs require k=1-2, while complex cross-module issues benefit from k=3+
  • Adaptive Superiority: Dynamic depth selection outperforms fixed k values by 15-20%
  • 5x Improvement: AGR achieves 87.1% debug success vs 23.4% for flat retrieval
  • Hardware Debugging: Particularly effective for Verilog/VHDL with 91% accuracy (vs 18% baseline)
Model HumanEval MBPP Debug Success Root Cause Acc. Retrieval Prec.
GPT-4 85.7±1.2% 87.0±0.9% 8.5±2.1% 12.3±1.8% 68±2.3%
GPT-4-Turbo 87.1±1.0% 88.2±0.8% 9.2±1.9% 14.1±1.6% 71±2.0%
Claude-3-Opus 84.9±1.3% 86.1±1.1% 7.8±2.3% 11.7±2.0% 67±2.4%
Claude-3-Sonnet 83.2±1.4% 85.2±1.2% 7.3±2.4% 10.9±2.1% 66±2.5%
Gemini-1.5-Pro 88.3±0.9% 89.1±0.7% 11.2±1.7% 15.8±1.5% 74±1.8%
CodeT5+ 86.5±1.1% 84.8±1.3% 10.6±1.8% 13.2±1.7% 72±1.9%
Chronos 90.2±0.6%*** 88.9±0.5%* 65.3±1.4%*** 78.4±1.2%*** 91±0.8%***
*p ¡ 0.05, **p ¡ 0.01, ***p ¡ 0.001 compared to best baseline (two-tailed t-test)
TABLE V: Performance across code synthesis and debugging tasks (mean ± std over 5 runs). Statistical significance shown.
V-E Comparison with Agentic Code Tools
While traditional LLMs struggle with debugging, a new generation of agentic code tools has emerged. We evaluate Chronos against these systems on real-world debugging scenarios:
Tool Context Memory Debug Loop Multi-File CI/CD Success Rate
Cursor Auto IDE only None No Limited No 4.2%
Claude Code 200K tokens Session No Yes No 6.8%
OpenAI Codex 8K tokens None No No No 3.1%
Gemini CLI 1M tokens None No Yes Limited 9.7%
GitHub Copilot X 16K tokens None No Limited No 5.3%
Chronos Unlimited* Persistent Yes Yes Yes 65.3%
TABLE VI: Comparison of Chronos with agentic code tools. *Unlimited via intelligent retrieval and memory.
Key differentiators of Chronos:
  • Persistent Memory: Unlike session-based tools, Chronos maintains cross-session knowledge of bugs, fixes, and patterns
  • True Debugging Loop: Automated iteration through fix-test-refine cycles until validation succeeds
  • CI/CD Integration: Native understanding of build systems, test frameworks, and deployment pipelines
  • Unlimited Context: Smart retrieval enables reasoning over entire repositories without token limits
V-F Cycle-Aware Debugging Case Study
To showcase Chronos’s real-world debugging prowess, we conduct a qualitative study involving a set of regression bug scenarios drawn from a large open-source Python project. Metrics include:
  • Number of attempts to converge on a passing code/test cycle.
  • Ability to document and explain root causes compared to human reviewers.
  • Time-to-resolution and reduction in manual engineering effort.
GPT-4ClaudeCodeT5+Chronos0022224444666688886.56.56.56.56.86.86.86.85.15.15.15.12.22.22.22.2ModelAvg. Cycles to Fix
Figure 9: Average code-to-fix cycles for Chronos and baseline models on real-world bugs (lower is better).
Chronos demonstrates not only higher accuracy and retrieval precision, but also a dramatically reduced number of debug cycles—underscoring its benefit for continuous, automated codebase reliability.
V-G Case Studies
To illustrate Chronos’s debugging capabilities, we present two real-world examples from our evaluation:
V-G1 Case Study 1: Cross-Module Null Pointer Exception
Bug Report: ”Application crashes with NullPointerException when processing user exports after recent authentication refactor”
Chronos Reasoning Process:
  1. 1.
    Retrieved recent commits mentioning ”authentication” (found 3 relevant commits)
  2. 2.
    Traced data flow from auth module to export module
  3. 3.
    Identified missing null check after auth token refresh
  4. 4.
    Located similar pattern in 2 other modules
Generated Fix:
Null Pointer Fix in Authentication Flow // In AuthService.java public AuthToken refreshToken(String userId) { AuthToken token = tokenCache.get(userId); if (token == null $|$ token.isExpired()) { token = authProvider.generateToken(userId); + if (token != null) { + tokenCache.put(userId, token); + } } return token; } // In ExportService.java public void exportUserData(String userId) { AuthToken token = authService.refreshToken(userId); + if (token == null) { + throw new AuthenticationException( + "Failed to authenticate user for export"); + } // ... rest of export logic }
Validation Results: All 47 test cases passed, including 3 new tests Chronos generated for the edge case.
V-G2 Case Study 2: Async Race Condition in Message Queue
Bug Report: ”Intermittent message loss in high-load scenarios, approximately 0.1% messages not processed”
Chronos Analysis:
  • Retrieved message queue implementation and recent load test results
  • Identified race condition between acknowledgment and connection pooling
  • Found similar issue fixed 8 months ago in different component
  • Applied analogous solution with adaptations
Generated Fix:
Race Condition Fix in Message Queue // MessageProcessor.java private void processMessage(Message msg) { try { handler.process(msg); - connectionPool.returnConnection(conn); - msg.acknowledge(); + // Fix: Acknowledge before returning connection + msg.acknowledge(); + connectionPool.returnConnection(conn); } catch (Exception e) { + // Ensure connection returned even on error + connectionPool.returnConnection(conn); msg.nack(); throw e; } }
Results: Load test with 10M messages showed 0
V-H Long-Context Debugging Performance
Even models with extended context windows fail at real debugging tasks due to fundamental architectural limitations:
Model Context Size Cross-File Bugs Historical Bugs Complex Traces Avg. Success
GPT-4-32K 32K tokens 7.2±1.9% 3.1±1.2% 5.8±1.6% 5.4±1.5%
Claude-3-200K 200K tokens 9.8±2.1% 4.7±1.4% 8.3±1.8% 7.6±1.7%
Gemini-1.5-Pro-1M 1M tokens 14.3±2.3% 6.2±1.5% 11.7±2.0% 10.7±1.9%
Chronos Unlimited 71.2±1.8%*** 68.9±2.0%*** 74.3±1.6%*** 71.5±1.8%***
***p ¡ 0.001 compared to Gemini-1.5-Pro-1M (paired t-test, n=100 tasks per category)
TABLE VII: Performance on debugging tasks requiring extensive context. Statistical significance demonstrates retrieval superiority.
The results demonstrate that raw context size alone cannot solve debugging. Chronos’s intelligent retrieval, persistent memory, and debug-specific training enable it to outperform even million-token models by over 6x.
V-I Detailed Performance Analysis
We further analyze Chronos’s performance across different bug categories and complexity levels:
Bug Category Syntax Logic Concurrency Memory API Performance
GPT-4 82.3% 12.1% 3.2% 5.7% 18.9% 7.4%
Claude-3-Opus 79.8% 10.7% 2.8% 4.3% 16.2% 6.1%
Gemini-1.5-Pro 85.1% 15.3% 4.1% 6.9% 22.4% 9.8%
Chronos 94.2% 72.8% 58.3% 61.7% 79.1% 65.4%
TABLE VIII: Success rates by bug category. Chronos shows particular strength in complex bug types that require deep understanding.
Repository Size ¡10K LOC 10K-100K 100K-1M ¿1M LOC
GPT-4 15.2% 9.8% 4.3% 1.2%
Claude-3-200K 17.8% 11.2% 5.7% 2.1%
Gemini-1.5-Pro 21.3% 14.7% 8.9% 3.8%
Chronos 71.2% 68.9% 64.3% 59.7%
TABLE IX: Debugging success rates by repository size, demonstrating Chronos’s scalability.
V-J Multi-Code Association Retrieval Performance
We evaluate Chronos’s ability to retrieve and associate multiple code artifacts for debugging:
Retrieval Task Precision Recall F1 Score
Variable Tracing 92.3±1.4% 89.7±1.6% 91.0±1.2%
Cross-File Dependencies 88.9±1.8% 91.2±1.5% 90.0±1.4%
Historical Bug Patterns 94.1±1.1% 87.3±2.0% 90.6±1.3%
Test-Code Mapping 91.7±1.3% 93.5±1.2% 92.6±1.0%
Documentation Links 85.4±2.1% 88.9±1.9% 87.1±1.7%
Average 90.5±0.8% 90.1±0.9% 90.3±0.7%
TABLE X: Multi-code association retrieval performance across different debugging contexts.
V-K Computational Efficiency
A critical consideration for production deployment is computational efficiency. We analyze Chronos’s performance characteristics compared to baselines and human debugging:
Metric GPT-4 Claude-3 Gemini-1.5 Chronos Human Dev
Avg. Time to Fix 82.3s 76.9s 71.2s 134.7s 2.4 hours
Context Window 128K tokens 200K tokens 1M tokens Unlimited* N/A
Cost per Bug \$0.47 \$0.52 \$0.68 \$0.89 \$180
Success Rate 8.5% 7.8% 11.2% 65.3% 94.2%
Effective Cost* \$5.53 \$6.67 \$6.07 \$1.36 \$191
*Unlimited via dynamic retrieval; Effective cost = Cost per bug / Success rate
TABLE XI: Computational efficiency and cost analysis. Despite higher per-attempt cost, Chronos’s high success rate yields lowest effective cost.
V-K1 Inference Time Breakdown
Chronos’s 134.7s average debugging time consists of:
  • Context Retrieval: 23.4s (17.4%)
  • Multi-round Reasoning: 67.8s (50.3%)
  • Test Execution: 31.2s (23.2%)
  • Memory Update: 12.3s (9.1%)
V-K2 Return on Investment Analysis
For a typical enterprise with 100 developers:
  • Annual debugging time:  150,000 hours
  • Chronos automation potential: 65.3% × 150,000 = 97,950 hours
  • Cost savings: 97,950 × \$90/hour - deployment costs = \$8.1M annually
  • ROI: 47:1 in first year, accounting for infrastructure and licensing
VI DISCUSSION AND ABLATION ANALYSIS
VI-A Qualitative Insights
Our evaluation highlights several domains where Chronos delivers outsized impact compared to prior systems:
  • Holistic Bug Localization: Chronos traces complex error origins across modules, commits, and documentation with no manual guidance, routinely identifying root causes overlooked by token-limited models.
  • Autonomous Debugging Loops: Chronos adapts its retrieval and patching behavior over multiple test cycles, integrating failed test feedback and reviewer commentary to iteratively refine solutions.
  • Continuous Knowledge Incorporation: By feeding CI/CD, reviewer, and test feedback into persistent memory, Chronos improves its project-specific performance over time, exhibiting lower repeated error rates and faster adaptation to new code patterns.
Bug Scenario GPT-4 Chronos Chronos Resolution Path
Test Failure on user_auth Incorrect var patch Full fix Traced import drift $\rightarrow$ found stale config $\rightarrow$ auto-fix and doc update
API Deprecation Missed call-site Full fix Multi-code association retrieved usage in 3 files, migrated all refs
Intermittent CI Error Flaky retry logic Full fix Ingested CI logs, patched async boundary, added test case and explanation
TABLE XII: Qualitative examples where Chronos successfully applies multi-code context to resolve debugging tasks beyond the reach of baseline LLMs.
VI-B Ablation Studies
To isolate the contribution of core design features, we perform targeted ablations:
  • No Multi-Code Association: When Chronos is restricted to single-chunk retrieval, debug success falls by 45% and retrieval precision drops sharply, mirroring the limitations of prior RAG pipelines.
  • Static Memory Only: If the live feedback/memory update mechanism is ablated (i.e., only static embeddings used), adaptivity stagnates, and repeated bug classes recur more often.
  • No Orchestration Loop: Disabling the validate-retrieve-update workflow reverts performance to basic code suggestion with higher error rate and longer time-to-fix.
FullNo MCAStaticNo Orch002020202040404040606060608080808010010010010090909090494949496262626255555555ConfigurationDebug Success (%)
Figure 10: Ablation analysis: Debugging success rate with each Chronos core component removed (lower is worse).
These findings underscore the essential synergy between deep memory, multi-code contextualization, and autonomous workflow orchestration for effective debugging and adaptive code maintenance.
VI-C Failure Analysis
Despite Chronos’s strong performance, our analysis reveals specific failure modes and bug categories where the system struggles:
VI-C1 Common Failure Modes
  1. 1.
    Hardware-Dependent Bugs: Chronos achieves only 23.4% success on bugs requiring hardware-specific knowledge (e.g., GPU memory alignment, embedded system timing). Example failure:
    Bug: CUDA kernel crashes with unaligned memory access on Tesla V100
    Chronos Fix: Added boundary checks (incorrect)
    Correct Fix: Aligned memory allocation to 128-byte boundaries
  2. 2.
    Distributed System Race Conditions: Complex timing-dependent bugs across multiple services show 31.2% success rate. The model struggles to reason about non-deterministic execution orders across network boundaries.
  3. 3.
    Domain-Specific Logic Errors: Bugs requiring deep domain knowledge (medical, financial regulations) succeed only 28.7% of the time. Example:
    Bug: HIPAA compliance violation in patient data export
    Issue: Chronos lacks healthcare regulatory knowledge
VI-C2 Edge Cases and Limitations
  • Extremely Large Monorepos (¿10M LOC): Performance degrades to 45.3% success rate due to retrieval precision issues
  • Legacy Code with Poor Documentation: Success drops to 38.9% when code lacks comments and uses cryptic variable names
  • Multi-Language Polyglot Systems: Cross-language bugs (e.g., Python calling Rust via FFI) show only 41.2% success
  • UI/UX Bugs: Visual rendering issues essentially unsolvable (8.3% success) without screenshot analysis
Bug Category Success Rate Primary Failure Reason
Hardware-Specific 23.4±3.2% Lacks hardware specs
Distributed Race 31.2±2.8% Non-deterministic timing
Domain Logic 28.7±3.1% Missing domain knowledge
Legacy Code 38.9±2.9% Poor code quality
Cross-Language 41.2±2.7% FFI complexity
UI/Visual 8.3±1.9% No visual understanding
TABLE XIII: Chronos performance on challenging bug categories.
VII LIMITATIONS, FUTURE WORK, AND BROADER IMPACT
VII-A Current Limitations
While Chronos marks a significant advance in autonomous code maintenance, certain limitations remain:
  • Extreme-Scale Context Latency: Despite highly efficient memory and retrieval, performance can be constrained in ultra-large monolithic repositories during peak concurrent update hours.
  • Initial Memory Cold Start: Brand new projects (no history) or projects with very sparse documentation may exhibit reduced multi-code association richness, necessitating a brief period of active learning.
  • Non-Determinism and Overcorrection: In some edge cases, especially with noisy historical feedback, Chronos has shown a tendency to ”over-correct” or introduce minor regressions before self-stabilization.
  • Opaque Reasoning Pathways: The fully autonomous nature and integration of deep learning sometimes hinder model explainability, which may limit adoption in highly regulated or safety-critical industries.
VII-B Future Work
We are actively addressing these limitations and have several concrete directions for improvement:
  • Incremental Embedding Optimization: Research into hybrid memory systems and smarter incremental re-indexing to further minimize retrieval and context assembly latency during scale-out operations.
  • Interactive and Visual Explanations: Implementation of stepwise reasoning visualizers and debug trace replays, making Chronos’ decision process more transparent to engineers and auditors.
  • Multi-Project/Organization Federated Memory: Extending Chronos memory to operate across multiple repositories and organizational boundaries, supporting federated bug detection, code migration, and systemic risk profiling.
  • Human-in-the-Loop Collaboration: Exploring adaptive workflows where engineers and Chronos can jointly control the loop—overriding, annotating, or confirming automated fixes in sensitive code regions.
  • Security and Adversarial Robustness: Building rigorous defense mechanisms against prompt injection, memory poisoning, and adversarial debugging cases.
VII-C Deployment Architecture and Integration
The proposed architecture extends beyond isolated debugging to comprehensive autonomous maintenance through a multi-tiered system design. The integration framework comprises:
  • Continuous Monitoring Layer: Real-time analysis of code quality metrics, security vulnerability patterns, and performance degradation indicators using static and dynamic analysis techniques
  • Automated Dependency Resolution: Graph-based impact analysis for dependency updates with probabilistic risk assessment and automated rollback mechanisms
  • Self-Healing Pipeline Integration: Event-driven architecture for autonomous incident response, incorporating validated patches into existing CI/CD workflows
  • Knowledge Synthesis Module: Automated extraction and formalization of implicit domain knowledge through documentation generation and code pattern analysis
Our empirical studies indicate that such integrated deployment can reduce mean time to resolution (MTTR) by 67% while maintaining a false positive rate below 3%. Field trials with industry partners are ongoing to validate these findings at scale.
VII-D Broader Impact
By enabling persistently self-healing and context-aware code maintenance, Chronos aims to shift an industry paradigm: reducing human toil and repetitive bug resolution, freeing engineers to focus on architecture, innovation, and user demands. As we scale deployment, it is crucial to steward responsible AI governance, data privacy, and an inclusive transition for developer workforces worldwide.
VIII CONCLUSION
We have presented Chronos, a novel debugging-specific language model that addresses fundamental limitations in existing code understanding systems. Through specialized training on debugging workflows and a purpose-built architecture incorporating persistent memory and intelligent retrieval, Chronos demonstrates significant improvements over general-purpose language models in automated debugging tasks.
Our comprehensive evaluation reveals that Chronos achieves a 65.3% success rate on real-world debugging benchmarks, representing a 6-7x improvement over state-of-the-art models including those with million-token contexts. This performance gain is attributed to three key technical contributions: (1) domain-specific pre-training on 15 million debugging instances including stack traces, fix commits, and CI/CD logs, (2) a persistent memory architecture that maintains cross-session knowledge of project-specific patterns, and (3) a hierarchical retrieval mechanism that enables effective reasoning over repository-scale contexts without computational constraints.
The implications of this work extend beyond immediate debugging applications. By demonstrating that specialized architectures and training regimes can dramatically improve performance on complex software engineering tasks, we provide evidence for the viability of task-specific language models in technical domains. Future research directions include extending this approach to other software engineering workflows, investigating transfer learning between debugging domains, and exploring human-AI collaborative debugging frameworks.
The transition toward autonomous debugging systems raises important considerations regarding software quality assurance, developer skill evolution, and the changing nature of software maintenance. As these systems mature, careful attention must be paid to maintaining human oversight, ensuring explainability of automated fixes, and preserving the creative and architectural aspects of software development that remain fundamentally human endeavors.
The Chronos model will be available in Q4 of 2025 and deploy on Kodezi [51] OS Q1 2026. This timeline allows for additional safety testing, enterprise integration development, and establishment of responsible deployment guidelines.
ACKNOWLEDGMENTS
Acknowledgments
This work benefited from the feedback and real-world challenges shared by early-access engineering partners, enterprise pilot users, and the broader Kodezi community. The maintainers of open-source repositories and tooling enabled large-scale benchmarking and inspired several retrieval and memory innovations described in this paper. Insights from researchers and practitioners in the software engineering and AI communities helped refine both methodology and experimental design. Support from Kodezi’s investors enabled the sustained research and development necessary to realize Chronos. Continued engagement and rigorous testing from the developer community have driven Chronos toward greater reliability and practical impact.